# Hofstadter-Toda spectral duality and quantum groups

### v1

@@ Conclusions
Concluding, we discussed and clarified the spectral relationship between the Hofstadter model in condensed matter physics and the relativistic Toda lattice in high-energy physics found by Hatsuda, Katsura, and Tachikawa [HKT16] in the framework of the representation theory of the elementary quantum group. Furthermore, we derived a formula parametrizing the energy spectrum of the Hofstadter model in the Brillouin zone in terms of elementary symmetric polynomials and Chebyshev polynomials, building on previous work on the Hofstadter model by Wiegmann and Zabrodin [WZ94] and on tridiagonal matrices by Molinari [Mol97]. We hope that our work will serve as a basis for a deeper understanding of the self-similarity properties of the Hofstadter butterfly and contribute to shed light on the connection between the Hofstadter and the relativistic Toda lattice models and, more generally, on the connection between noncommutative quantum geometry and the quantum world.
@@ Acknowledgements
P. M. is supported by the Japan Science and Technology Agency (JST) of the Ministry of Education, Culture, Sports, Science and Technology (MEXT), JST CREST Grant No. JPMJCR19T2, the Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Early-Career Scientists KAKENHI Grant No. JP23K13028 and No. JP20K14375. V. P. is supported by the JST CREST Grant. No. JPMJCR19T2. X. S. is partially supported by KAKENHI Grant No. 21K03259 and Grants from Joint Project of OIST, Hikami Unit and University of Tokyo.
The authors certify that he has no affiliations with or involvement in any organization or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript.

### v2

@@ Conclusions
Concluding, we discussed and clarified the spectral relationship between the Hofstadter model in condensed matter physics and the relativistic Toda lattice in high-energy physics found by Hatsuda, Katsura, and TachikawaHatsuda, Katsura, and Tachikawa (2016) in the framework of the representation theory of the elementary quantum group. Furthermore, we derived a formula parametrizing the energy spectrum of the Hofstadter model in the Brillouin zone in terms of elementary symmetric polynomials and Chebyshev polynomials, building on previous work on the Hofstadter model by Wiegmann and ZabrodinWiegmann and Zabrodin (1994) and on tridiagonal matrices by MolinariMolinari (1997). We hope that our work will serve as a basis for a deeper understanding of the self-similarity properties of the Hofstadter butterfly and contribute to shed light on the connection between the Hofstadter and the relativistic Toda lattice models and, more generally, on the connection between noncommutative quantum geometry and the quantum world.
@@ Acknowledgements
P. M. thanks Hosho Katsura for useful discussions. P. M. is supported by the Japan Science and Technology Agency (JST) of the Ministry of Education, Culture, Sports, Science and Technology (MEXT), CREST Grant No. JPMJCR19T2, the Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Early-Career Scientists KAKENHI Grants No. JP23K13028 and No. JP20K14375. V. P. is supported by the JST CREST Grant. No. JPMJCR19T2 and by Marie Skłodowska-Curie Individual Fellowship (project number 101063362). X. S. is partially supported by KAKENHI Grant No. JP21K03259 and Grants from Joint Project of OIST, Hikami Unit and the University of Tokyo.
Data Availability Statement